If demand function is given as the following: 

Qz = 230 -2.75 Pz + 0.5 I + 1.2 Pm + 0.6A 

Where Qz is quantity of Good z sold, Pz is price of Good z per unit, I is per capita income, Pm is price of competitor and A is the amount of advertising spent. 

Current values:  Pz= RM 55 I= RM 9000    Pm= RM 50     A =RM 12,000

a) Calculate price elasticity of demand for Good z.



1
Expert's answer
2021-05-04T07:38:32-0400
"Q_z=230-2.75\times55+0.5\times9000+1.2\times50+0.6\times12000=11838.75"

"E=-2.75\frac{55}{11838.75}=-0.013"